5.) A 14 ml solution of KOH has a concentration of 0.315 M. Calculate the pH and then calculate the amount of KOH in grams in solution.
1
Expert's answer
2020-05-07T14:08:42-0400

KOH = K+ + OH-

[OH-] = 0.315 M

pOH = - log[OH-]

pOH = 0.502

pH = 14 - 0.502 = 13.5


(14 / 1000) L * 0.315 M = 0.00441 mol

M (KOH) = 56.11 g/mol

m = 56.11 * 0.00441 = 0.247 g